### v1

# Learning More by Seeing Less: Line Drawing Pretraining for Efficient, Transferable, and Human-Aligned Vision

Abstract
Despite remarkable progress in computer vision, modern recognition systems remain limited by their dependence on rich, redundant visual inputs. In contrast, humans can effortlessly understand sparse, minimal representations like line drawings—suggesting that structure, rather than appearance, underlies efficient visual understanding. In this work, we propose using line drawings as a structure-first pretraining modality to induce more compact and generalizable visual representations. We show that models pretrained on line drawings develop stronger shape bias, more focused attention, and greater data efficiency across classification, detection, and segmentation tasks. Notably, these models also exhibit lower intrinsic dimensionality, requiring significantly fewer principal components to capture representational variance—echoing the similar observation in low dimensional efficient representation in the brain. Beyond performance improvements, line drawing pretraining produces more compressible representations, enabling better distillation into lightweight student models. Students distilled from line-pretrained teachers consistently outperform those trained from color-supervised teachers, highlighting the benefits of structurally compact knowledge. Finally, we demonstrate that the pretraining with line-drawing can also be extended to unsupervised setting via our proposed method ”learning to draw”. Together, our results support the view that structure-first visual learning fosters efficiency, generalization, and human-aligned inductive biases—offering a simple yet powerful strategy for building more robust and adaptable vision systems.
Introduction
Despite remarkable progress in computer vision, modern recognition systems remain fundamentally constrained by their reliance on dense, high-fidelity visual inputs. State-of-the-art models, particularly those based on deep convolutional and transformer architectures, have demonstrated impressive accuracy across a variety of benchmarks. However, their success is heavily contingent on exposure to vast datasets filled with richly textured, color photographic imagery. This dependency introduces significant inefficiencies—both in terms of data requirements and generalization capabilities—limiting the robustness of these systems when confronted with abstract or atypical visual inputs (Geirhos et al. 2019; Schlegel et al. 2015).
Figure 1: Pretraining with Line Drawing Leads to efficient, transferable and Human-like visual representation.
In contrast, humans can effortlessly interpret line drawings—minimal, structure-preserving representations that strip away redundant surface features. This ability suggests that the human visual system is tuned to extract and reason over efficient structural information such as shape and topology, rather than relying on unstable texture or color. Neuroscientific fMRI studies support this view, showing that line drawings elicit robust, category-specific activity patterns throughout the visual cortex—comparable to those induced by full-color photographs (Walther et al. 2011). Similarly, Shin et al. (2008) also demonstrates that abstract visual symbols like icons also elicits similar neuronal activation in the brain as their corresponding natural images.
Psychological evidence further supports the transformative power of minimal representations in visual learning. Studies reveal that drawing practice enhances perceptual skills and object recognition in humans (Kozbelt 2001; Chamberlain et al. 2019; Perdreau and Cavanagh 2015). Artists develop specialized visual cognition through prolonged exposure to structural information—improved visual memory, enhanced structural analysis, and refined attention to diagnostic features (Perdreau and Cavanagh 2015; Schlegel et al. 2015). These enhancements are believed to be associated with minimal representations learning (Perdreau and Cavanagh 2014)—by eliminating distracting surface features, it compells the visual system to extract and encode deeper and more robust structural relationships.
Motivated by these insights, we propose a structure-first visual learning paradigm centered on line drawing pretraining for better representation learning. Our central hypothesis is that by training vision models to first understand the world through minimal structural cues, we aim to induce representations that are more compact, efficient, generalizable, and human-aligned. Thanks to the recent development from Chan, Durand, and Isola (2022) , we now can accurately convert photographic color images to its minimal line-drawing form automatically at scale. Our results show that models pretrained on line drawings exhibit stronger shape bias, more focused attention, and lower intrinsic dimensionality—mirroring characteristic findings of human minds.
These models also yield substantial gains in data efficiency and distillation, outperforming color-supervised teachers in student performance even when matched on classification accuracy. Additionally, we show that Line drawing pretraining can also be extended to unsupervised pretraining with similar performance gain via a ”learning to drawing” strategy, effectively training the model to perform drawing exercise in order to train the visual perception.
Together, our findings suggest that minimal, structure-centered pretraining is a powerful strategy for fostering robust, efficient, and transferable vision systems—grounded in the computational principles that underlie human perception.
Related Work
Line Drawing Synthesis and Structural Representation   Recent advances in line drawing generation have focused on balancing geometric fidelity and semantic preservation. The work by (Chan, Durand, and Isola 2022) pioneered an unsupervised approach using geometry and semantic losses to translate photographs into sparse yet informative sketches, circumventing the need for paired datasets. Earlier methods relied on supervised training with stroke annotations or 3D geometry, limiting their applicability to arbitrary photographs (Das et al. 2020; Ganin et al. 2018; Ha and Eck 2018; Smirnov et al. 2020). Concurrent work in unpaired translation, such as CycleGAN (Zhu et al. 2017) and UPDG (Yi et al. 2019), achieved domain adaptation but struggled with structural coherence in sparse outputs. Our framework builds on these foundations by reorienting line drawing synthesis as a pretraining objective, prioritizing shape encoding over pixel-level reconstruction.
Self-Supervised Learning Paradigms   Self-supervised pretraining has emerged as a powerful tool for learning robust visual representations. Masked autoencoders (He et al. 2021) demonstrated that reconstructing masked image patches fosters feature generalizability, while DINO (Caron et al. 2021) showed that self-distillation without labels enables Vision Transformers to learn semantic segmentation properties and object boundaries without explicit supervision. However, these methods either prioritize low-level pixel reconstruction or rely on complex view-matching objectives rather than explicitly modeling structural understanding. Our ”learning to draw” approach diverges by framing line drawing synthesis as a psychologically grounded pretext task, mirroring human skill acquisition through iterative shape abstraction. This aligns with evidence from cognitive science, where observational drawing practice enhances perceptual discrimination and neural reorganization.
Methods to Improve Shape Bias   The texture bias of convolutional neural networks has motivated efforts to instill shape-centric reasoning. Work by (Geirhos et al. 2019) showed that stylized training data improves shape sensitivity, but this did not scale well. Other approaches like (Hermann, Chen, and Kornblith 2020) and (Lee et al. 2022) respectively used human-like and shape-guided data augmentation techniques to induce more shape bias. Work by (Dapello et al. 2023) focused on architectural improvements and induced shape bias through sparsity constraints in transformer architectures. Our work bridges these insights by using line drawings-natural sparse representations-to explicitly steer models toward geometric regularization. Unlike augmentation strategies, our method leverages the inherent spatial regularity of line art to improve cross-domain generalization.
Sparsity in Neural Representations   Sparse coding has long been theorized as a mechanism for efficient visual processing. Recent work (Dapello et al. 2023) demonstrates that Top-K activation sparsity enhances shape bias by emphasizing part-based representations, while masked autoencoders (He et al. 2021) use spatial sparsity to improve feature learning. Line drawings naturally instantiate this principle, distilling scenes to their essential contours. Our method operationalizes sparsity through the unsupervised drawing task, encouraging models to discard texture noise and focus on structural invariants — a process analogous to artists’ mental decomposition of complex forms (Perdreau and Cavanagh 2014).
Method
Supervised Pretraining
We employ a two-stage supervised pretraining approach using the STL-10 dataset with stratified sampling across data subsets ranging from 5% to 100% of the full training corpus. Our methodology compares transfer learning approaches against a baseline control where models train directly on natural color images without pretraining, establishing the performance ceiling for direct supervised learning.
We convert the original STL-10 (Coates, Lee, and Ng 2011) training images to line drawings using the method from (Chan, Durand, and Isola 2022), which extracts the most crucial features in line form. For the artistic style variants, we apply style transfer techniques from (Huang and Belongie 2017) to generate stylized versions of the original images, creating distinct visual abstractions that test semantic transfer across diverse visual domains.
Figure 2: Different augmentations of original image
In the transfer learning protocol, we first pretrain on source domains including line drawings and various artistic styles, all using identical class labels as the target domain. Line drawings serve as our primary focus, providing structural information without color, while the artistic styles function as arbitrary augmentations to compare against. We then fine-tune the pretrained models on natural color images, allowing them to leverage learned semantic representations while adapting to target visual characteristics.
The training configuration employs adaptive learning rates (0.001-0.01) and batch sizes (4-32) that scale with dataset size, along with early stopping (patience=10). Both training stages utilize identical data augmentation strategies including random resized cropping, horizontal flipping, and rotation. Models are evaluated the STL10 test sets using center-cropped images, with comprehensive testing across all data percentage levels under fixed random seeds for reproducibility.
Learning to Draw Architecture
We propose a novel approach to visual representation learning inspired by the cognitive processes underlying artistic drawing. Our method, Learning to Draw, trains a model to translate photographs into line drawings using an unsupervised objective that mirrors human perceptual development—where drawing enhances structural understanding by emphasizing geometry over texture.
Our architecture adopts an asymmetric encoder-decoder design, conceptually similar to Masked Autoencoders (He et al. 2021), but with a different goal: instead of reconstructing masked patches, we distill essential structural components from images. The encoder is a modified ResNet-18 with multi-scale feature fusion across layers 2–4, aligned and concatenated to capture both local and global structure. We reduce the first convolution stride to 1 and use a 3×3 kernel to retain fine details, followed by a 1×1 convolution for channel adaptation.
The decoder, following (Chan, Durand, and Isola 2022), is intentionally lightweight: two transposed convolutions and a final 7×7 convolution with sigmoid activation. This asymmetric setup creates an information bottleneck that encourages the encoder to extract compact, structure-focused representations.
Training is performed on unpaired data from 10k COCO images and 14,224 Anime Colorization sketches. We use the losses from (Chan, Durand, and Isola 2022), combining adversarial, geometric (via monocular depth), semantic (via CLIP features), and cycle-consistency losses. The model is trained for 200 epochs with batch size 8, using Adam optimizer (LR=0.0002, with linear decay starting at epoch 100).
Results
Pretraining with Line Drawings
The results presented in Table 1 reveal important insights into the effects of training strategy on model performance for classification tasks involving both color images and line drawings. Initializing training with line drawings (STL10-Line) followed by color images (STL10) leads to an improvement in generalization to line drawings, with accuracy rising from 16.67% to 29.53%. At the same time, color image classification accuracy also increases from 68.79% to 73.58%, indicating that structural features learned from line drawings provide a beneficial inductive bias for learning color representations.
Training Strategy Color Acc Line Acc
Color Only $68.79$ $16.67$
Line Only $10.67$ $62.62$
Line→Color $73.58$ $29.53$
Color→Line $13.43$ $66.92$
Table 1: Classification accuracy (%) for different training strategies.
The experiments also demonstrate a pronounced asymmetry in performance depending on the order of training modalities. When models are trained first on color images and then on line drawings, color image classification accuracy drops sharply to 13.43%, despite strong performance on line drawings (66.92%). Conversely, training first on line drawings and then on color images results in high color image accuracy (73.58%) and moderately line drawing accuracy (29.53%). This suggests that features derived from line drawings are more generalizable and facilitate adaptation to color images, whereas color features are more specific and less transferable to line drawings.
These findings have practical implications for the design of image learning systems. Line-first training offers a robust initialization that is particularly effective when computational resources are limited, while simultaneous training is optimal for maximizing performance. The results also highlight the risk of catastrophic forgetting when adapting models trained solely on color images to line drawings. Overall, the results support the notion that structural features learned from line drawings are more transferable and beneficial for generalization across different visual modalities, and suggest that future work should consider modality-specific architectural components and regularization techniques.
Training Strategy ImageNet-1k Accuracy
Color Only 67.865
Line → Color 68.823
Table 2: ImageNet-1k (Russakovsky et al. 2015) classification accuracy (%) for YOLO-v8 backbone (Reis et al. 2023) (11M parameters) under different training strategies. All experiments conducted at 224×224 resolution.
To validate the robustness of our findings beyond the ResNet architecture, we conducted additional experiments using a YOLO-v8 backbone with comparable model complexity (11M parameters). The training protocol consisted of 32 epochs of initial pretraining followed by 64 epochs of color image training, maintaining consistency with our sequential learning paradigm while adapting to the computational constraints of the larger-scale ImageNet-1k dataset.
The results presented in Table 2 demonstrate that the benefits of line drawing pretraining extend beyond ResNet architectures to modern detection-based backbones. The YOLO-v8 model pretrained on line drawings achieves 0.68823 accuracy on ImageNet-1k classification, representing a 0.958 percentage point improvement over the color-only baseline (0.67865). While the absolute magnitude of improvement is smaller than observed in our STL10 experiments (Table 1), the consistent direction of the effect across architectures and datasets provides strong evidence for the generalizability of our approach.
The reduced magnitude of improvement on ImageNet-1k compared to STL10 can be attributed to several factors. First, ImageNet-1k represents a substantially more challenging classification task with 1000 classes compared to STL10’s 10 classes, potentially limiting the relative impact of pretraining strategies. Second, the YOLO-v8 backbone incorporates architectural innovations designed for object detection, including feature pyramid networks and attention mechanisms that may already capture some of the structural biases induced by line drawing pretraining.
Notably, the YOLO-v8 results corroborate our central hypothesis that line drawing pretraining induces beneficial inductive biases that transfer across visual domains. The consistency of this effect across both discriminative (ResNet) and detection-oriented (YOLO) architectures suggests that the underlying mechanism—enhanced focus on structural features—is architecture-agnostic and represents a fundamental property of the pretraining data rather than specific architectural interactions.
These findings have important implications for practical deployment scenarios. The YOLO-v8 experiments demonstrate that line drawing pretraining can be effectively integrated into modern computer vision pipelines without requiring specialized architectural modifications. Furthermore, the maintained performance benefits on the large-scale ImageNet-1k benchmark indicate that our approach scales beyond smaller datasets and toy problems to real-world classification tasks.
The convergent evidence from both ResNet-18 (Table 1) and YOLO-v8 (Table 2) architectures establishes line drawing pretraining as a broadly applicable technique for improving visual representation learning, with potential applications spanning classification, detection, and other computer vision tasks.
Data Efficiency
Figure 3: Data efficiency results of pretraining with various augmentations (line drawings, Matisse style, Mondrian style, Picasso style) of STL10 (Coates, Ng, and Lee 2011) and finetuning on the STL10.
Our experimental results provide compelling evidence for the data efficiency advantages of line drawing pretraining over conventional color-only training approaches. As illustrated in Figure 3, the Line-Color pretraining method demonstrates superior learning efficiency across all training data regimes, achieving the baseline color accuracy (69%) with only 70% of the training data—representing a substantial 30% reduction in labeled data requirements. This efficiency gain manifests early in the training process, with the Line-Color approach consistently outperforming the color-only baseline by 3-5 percentage points at equivalent data volumes. Notably, the learning trajectory exhibits a steeper ascent during the initial training phases, suggesting that line drawing pretraining provides a more structured optimization landscape that facilitates faster convergence to target performance levels.
The comparative analysis with artistic style controls (Matisse-Color, Mondrian-Color, and Picasso-Color) further validates the specificity of line drawing advantages. While these alternative visual abstractions yield improvements over the color-only baseline, they are significantly less effective than line drawings in achieving data efficiency. The artistic style methods require substantially more training data to reach baseline performance compared to the line drawing approach. This differential performance strongly supports our theoretical framework that line drawings tap into fundamental shape and topology processing mechanisms inherent to visual recognition systems. The reduced effectiveness of alternative abstractions demonstrates that the efficiency gains are not merely artifacts of stylistic variation, but rather reflect the unique capacity of line drawings to provide structured, shape-preserving visual representations that align with the computational principles underlying human visual processing.
More Focused Attention
Figure 4: GradCAM visualizations comparing attention heat maps for models trained on color images only and models pretrained on line drawings then color images.
To understand the underlying mechanisms driving the aforementioned performance improvements observed, we analyzed the attention patterns and activation distributions of models trained under different pretraining strategies. Figure 4 presents GradCAM (Gildenblat and contributors 2021) visualizations comparing ResNet18 models trained exclusively on STL10 color images versus models pretrained on STL10-Line followed by STL10 color images. The visualizations reveal striking differences in attention allocation between the two training approaches.
Figure 5: Number of distinct high activation regions across all STL10 Test images. High activation regions are marked as the number of regions above the 85th percentile in GradCam activation.
Models pretrained on line drawings demonstrate markedly more focused attention patterns, with activation heat maps concentrated on the primary subject matter in each image. For the cat image (top row), the linedrawing-pretrained model exhibits concentrated attention on the cat’s facial features and body contours, while the color-only model shows more diffuse activation across background elements. Similarly, for the airplane image (bottom row), the line-pretrained model focuses attention primarily on the aircraft’s structural elements, whereas the color-only model displays scattered activation patterns across both the aircraft and background sky regions. To quantatitively confirm such representation difference, we isolate the highly activated region (red region in GradCAM) and compute the region count for each image. Across all testing cases in STL-10, we show that Line-drawing pretrianing leads to larger number of images with only one compact attention region while training with Color Only leads to more cases with two or more fragmented attention regions (Figure 5, see supplementary for detailed methods for region counting).
Figure 6: Neuron tuning curves of base ResNet18 vs ResNet18 with line drawing pretraining. (Normalized activations over 500 test images)
Sparse Neuronal Tuning Curves
The neuron tuning curve presented in Figure 6 provides quantitative support for these qualitative observations. We evaluate the activation tuning curve for the last layer of the trained ResNet-18. Each neuron in Figure 6 represents one channel. We take the average firing activations from each channel and normalize them according their maximum firing activation within each setting. The resulting activation tuning curve shows that ResNet18 pretrained on line drawings and then fine-tuned on natural images (blue curve) exhibits a much sharper tuning curve in activation compared to the standard counterpart (red curve). This pattern indicates that strong responses are concentrated among a small subset of neurons, while most units remain weakly or sparsely engaged. Such a rapid fall-off shows an increased sparsity and efficiency in neural representations.
Instead of distributing activation across many features, the line-pretrained network allocates its capacity to only the most informative neurons for the task, effectively silencing those features that are irrelevant or noisy. The base model, by contrast, shows a more gradual decrease, activating a broader collection of neurons, including many that likely encode non-discriminative or redundant information.
This selective focus echoes principles of efficient coding: the model learns to emphasize features that matter for object recognition while suppressing spurious or noisy activations. Notably, this shift toward more discriminative, quieter representations is not accompanied by any loss—and indeed comes with a gain—in performance. The line-pretrained model achieves stronger accuracy on both minimal and natural images.
These results encapsulate the representational transformation induced by minimal pretraining. Exposure to line drawings encourages the model to specialize its detectors for structural, geometric cues and to silence those features that do not contribute to meaningful classification. This focused efficiency visually and quantitatively captures why minimal visual experience builds more robust and transferable visual codes.
Increased Shape Bias
[…]
To further assess the inductive biases learned by our models, we evaluated shape versus texture preferences using the model-vs-human benchmark (Geirhos et al. 2021). This benchmark presents cue-conflict images in which shape and texture cues are intentionally mismatched, allowing for a direct comparison of classification strategies between models and human observers.
Figure 7 displays the distribution of shape and texture decisions across object categories (these are the 7 categories in the cue-conflict dataset that coincide with STL10 classes). Models pretrained on line drawings and subsequently fine-tuned on color images demonstrate a pronounced shift toward shape-based decisions, as indicated by the clustering of yellow circles toward higher fractions of shape choices. In contrast, models trained only on natural images show a greater reliance on texture cues, with more blue circles distributed toward the texture-dominant end of the spectrum. This pattern suggests that line drawing pretraining encourages the model to prioritize global structural features over local textural information when resolving ambiguous visual input.
These results indicate that line drawing pretraining not only enhances classification performance and attention focus, but also aligns model decision-making more closely with human perceptual strategies. This increased shape bias likely contributes to the improved robustness and generalization observed in downstream tasks, as models become less reliant on superficial texture cues and more attuned to the underlying structure of visual objects.
Benefits in Semantic Segmentation
We further show that pretraining on line drawings enhances transferability to dense prediction tasks such as semantic segmentation. To evaluate this, we fine-tune ResNet-18 models—initialized either from color-only training or from line-drawing → color training—on the ADE20K benchmark.
As shown in Figure 8, the model pretrained with line drawings consistently outperforms the color-only baseline across multiple semantic categories. The improvement is especially pronounced for object-centric classes (e.g., “chair”, “car”), where shape plays a dominant role in visual discrimination. This demonstrates that structural priors learned from minimal representations translate effectively to tasks requiring precise spatial reasoning.
These results extend our findings beyond classification and detection, highlighting that line drawing pretraining induces representations that generalize across both sparse and dense vision tasks.
Figure 8: Different line drawing pretraining methods on ResNet18 generalizing to segmentation on ADE20K.
Benefits in Object Detection
Here, we also run similar experiments in Figure 3 on object detection task using PASCAL VOC dataset. Figure 9 illustrates that models pretrained on line drawings consistently outperform color-only models across all training data regimes. Notably, our Linedrawing→Photo (Ours) model trained with just 90% of the data achieves an mAP of 56.74, surpassing the Photo Only model trained with the full 100% dataset, which reaches 56.14. This highlights a key advantage of our approach: better performance with less data. The performance gap is particularly prominent in the low-data regime—for example, at 10% training data, our model achieves 22.16 mAP compared to 18.17 for the color-only baseline. These findings suggest that pretraining on structure-focused line drawings enables models to extract more efficient and transferable features for object detection, leading to significant gains in data efficiency and generalization.
[…]
Compact and Transferable Representations
One of the core advantages of line-drawing pretraining is the emergence of compact and structurally focused internal representations. To investigate this, we first analyze the intrinsic dimensionality of neural activations in a ResNet-18 model trained under two regimes: color-only vs. line-drawing → color. We extract penultimate layer activations across STL10 test images and compute principal components (PCs) via PCA.
As shown in Figure 10, the line-drawing-pretrained model requires only 18 PCs to explain 90% of the variance, compared to 30 for the color-only model—suggesting that the model organizes information more economically and filters out redundant features. This compressed manifold structure aligns with principles of efficient coding and complements earlier findings on activation sparsity (Figure 6).
We further examine whether this structural compactness benefits model distillation. Using a standard teacher-student framework, we compare the performance of student models (MobileNetV1, ResNet-8, VGG-8) distilled from ResNet-18 teachers trained with either color-only or line-drawing → color supervision. As shown in Table 3, students distilled from line-pretrained teachers consistently outperform those trained from color-only teachers. For example, a MobileNetV1 distilled from a line-drawing-pretrained teacher (Pretrained with 70%-data to reduce the confounding factor of a stronger teacher) achieves 75.48% accuracy—exceeding the 73.83% achieved by the color-only teacher while the two teacher share almost the same object recognition performance.
These results suggest that line-drawing pretraining not only reduces the complexity of internal representations, but also produces knowledge that is easier to transfer, yielding more performant student models.
[…]
Teacher Teacher Acc Student Distillation Acc
[…]
Color (3) 68.89 MobileNetV1 73.83
[…]
Color (3) 68.89 ResNet-8 68.81
[…]
Color (3) 68.89 VGG-8 73.15
[…]
Benefits of Unsupervised Line Drawing Generation
Training Strategy Color Image Accuracy
Color Training Only 68.79%
[…]
Learning to Draw $\rightarrow$ Color Training 73.30%
Table 4: Comparison of unsupervised Learning to Draw pretraining on COCO and then finetuning on STL10 color image classification accuracy.
Our unsupervised Learning to Draw approach—training models to generate line drawings from color images without paired supervision—significantly improves visual recognition. The results presented in Table 4 demonstrate the effectiveness of our proposed unsupervised Learning to Draw approach as a pretraining strategy for visual recognition. It achieves 73.30% accuracy on STL10, surpassing color-only training by 4.51 points and reaches almost the same level as the supervised line-drawing pretraining. These results validate our hypothesis: learning to synthesize line drawings guides models to extract structural, shape-centric representations, even without explicit supervision.
Conclusion
Our findings show that line drawing pretraining induces compact, structure-aware representations that generalize across tasks and enable more effective distillation into lightweight models. This structure-first learning approach improves performance, enhances data efficiency, and yields representations that are easier to compress and transfer. More broadly, our results offer a computational lens on the longstanding observation that humans can easily understand line drawings—suggesting that prioritizing geometry over appearance may support more general and efficient visual reasoning, both in artificial and biological systems.